OpenStudy (anonymous):

What is the equation in point-slope form of the line passing through (4, 0) and (2, 6)? y = 4x − 2 y = 2x − 4 y = −3(x − 4) y = 3(x + 4)

OpenStudy (satan ):

Use the slope formula to find the slope first, what do you get?

OpenStudy (anonymous):

idk really

OpenStudy (anonymous):

im kinda retarded when it come to math

OpenStudy (satan ):

Slope formula: \(\Large m=\frac{y_{2}-y_{1}}{x_{2}-x_{1}}\)

OpenStudy (anonymous):

m=6-0 over 2-4

OpenStudy (anonymous):

6 over negative 2?

OpenStudy (satan ):

Yes, which simplifies to -3. :) Now, do you know how to use the point-slope formula?

OpenStudy (anonymous):

but there are 4 answers

OpenStudy (anonymous):

so its y=-3(x-4)?

OpenStudy (satan ):

By process of elimination, that's the only one with a slope of -3, so yea :P

OpenStudy (anonymous):

thank you soo much

OpenStudy (satan ):

You're weclome